# Atlas Gaussians Diffusion for 3D Generation with Infinite Number of Points

### v1

@@ Abstract
Using the latent diffusion model has proven effective in developing novel 3D generation techniques. To harness the latent diffusion model, a key challenge is designing a high-fidelity and efficient representation that links the latent space and the 3D space. In this paper, we introduce Atlas Gaussians, a novel representation for feed-forward native 3D generation. Atlas Gaussians represent a shape as the union of local patches, and each patch can decode 3D Gaussians. We parameterize a patch as a sequence of feature vectors and design a learnable function to decode 3D Gaussians from the feature vectors. In this process, we incorporate UV-based sampling, enabling the generation of a sufficiently large, and theoretically infinite, number of 3D Gaussian points. The large amount of 3D Gaussians enables high-quality details of generation results. Moreover, due to local awareness of the representation, the transformer-based decoding procedure operates on a patch level, ensuring efficiency. We train a variational autoencoder to learn the Atlas Gaussians representation, and then apply a latent diffusion model on its latent space for learning 3D Generation. Experiments show that our approach outperforms the prior arts of feed-forward native 3D generation.
Figure 1: We propose the Atlas Gaussians representation for 3D generation. Our method supports both unconditional (left) and conditional (middle) generation with great diversity. With Atlas Gaussians, we can generate a sufficiently large, and theoretically infinite, number of 3D Gaussian points. To demonstrate this, 100K Gaussian points generated by our method are shown (right). Additionally, we iteratively sample 32 patches of the shape, displaying each set in one of four subfigures (right).
@@ 1 Introduction
3D generation has become increasingly important across various domains, including virtual reality, gaming, and film production. Recent advances in diffusion models [32, 91] have improved the 3D generation quality, offering superior performance over previous methods, such as variational autoencoders (VAEs) [46, 68] and generative adversarial networks (GANs)[23, 21].
@@
Despite progress, the effectiveness of 3D generation still falls short when compared to 2D generation models [33, 79, 83]. The robust performance in 2D generation is largely due to the effective integration of VAEs with latent diffusion models (LDMs) [82]. A primary challenge hindering the complete success of this paradigm in 3D generation is the development of a high-fidelity 3D representation that can be efficiently embedded into a low-dimensional latent space. Pioneering efforts have applied diffusion models to various traditional 3D representations, such as point clouds [112, 120, 63], meshes [60], occupancy fields [118, 114] and signed distance functions [13, 50]. However, these approaches often focus solely on modeling geometry without considering the appearance attributes. More recently, a notable attempt [48] has designed a VAE that uses volume rendering techniques [66] to incorporate appearance modeling. Nevertheless, volume rendering presents inherent limitations, including slow rendering speeds and constrained rendering resolutions. To overcome these limitations, we have developed a VAE that leverages the latest 3D Gaussian representation [45], which significantly improves both the quality and speed of rendering.
@@
Designing a VAE based on 3D Gaussians presents considerable challenges. The first challenge lies in creating an efficient decoder capable of mapping low-dimensional latents to 3D Gaussians. Existing methods for 3D Gaussian decoding predominantly focus on reconstruction tasks and typically lack an information bottleneck design [97, 109], thus failing to provide low-dimensional latent directly. Alternatively, they often require multiple complex and interdependent components [105, 121]. The second challenge involves generating a sufficiently large number of 3D Gaussians efficiently, since high-quality rendering necessitates an adequate quantity of these Gaussians. Some current methods [105, 121] address this by employing additional complex point-upsampling networks to increase the number of 3D Gaussians, which inherently limits the number of Gaussians that can be generated. Other techniques [97, 109] utilize image representations to generate a large number of 3D Gaussians. However, for all these methods, more network parameters are usually required as the number of 3D Gaussians increases.
@@
To address the challenges of designing VAEs for 3D Gaussians, we propose Atlas Gaussians, a new representation for 3D generation. This representation is inspired by surface parameterization [20], a fundamental technique in many graphics applications where surface attributes are sampled and stored in a 2D texture map. Specifically, Atlas Gaussians model the shape as a union of local patches, with each patch decoding 3D Gaussians via UV-based sampling. By parameterizing 3D Gaussians in the UV space, we can easily generate a sufficiently large, and theoretically infinite, number of 3D Gaussians. Unlike traditional surface parameterization approaches, the UV mapping in Atlas Gaussians are learned end-to-end using a corner-encoding scheme. A significant advantage of Atlas Gaussians is that the sampling process does not require additional network parameters as the number of 3D Gaussians increases.
@@
We design a transformer-based decoder to map low-dimensional latents to Atlas Gaussians. This decoder achieves geometry and appearance disentanglement by design, facilitating faster convergence and improved representation capabilities. Leveraging the local awareness of Atlas Gaussians, we also reduce computational complexity by decomposing the self-attention layers. Finally, the latent space learned by our VAE can be applied to existing latent diffusion models efficiently.
@@
Note that in contrast to the major approach in 3D generation that uses the multi-view representation [100, 61, 59, 86], our approach is inherently 3D-based. Therefore, a key advantage is that we do not need to address the challenging multi-view consistency issue associated the multi-view representation. We take inspiration from the literature that 3D-based neural networks learn better 3D features for recognition and outperform multi-view based networks that leverage massive 2D labeled data. Moreover, the rendering module of Atlas Gaussians allows representation learning from images.
@@
• We design a new transformer-based decoder to map low-dimensional latents to Atlas Gaussians, achieving geometry and appearance disentanglement by design along with efficient computation.
@@ 2 Related work
3D representation. 3D reconstruction and generation benefit from different 3D representations by leveraging their unique properties. These representations include explicit representations [103, 67, 80, 112, 120, 63, 94, 104, 108, 95, 1, 18, 60, 71, 88, 25, 10, 45, 97, 109, 121, 96] and implicit representations [73, 11, 65, 50, 119, 87, 41, 38, 107, 66, 113, 6, 26, 8, 5, 69, 102]. In this paper, we focus on 3D Gaussians [45, 97, 109, 121, 96, 29], which possess high-quality rendering procedures that allow learning from image supervisions. However, existing results mainly focus on reconstructing 3D Gaussians. Our goal is to push the state-of-the-art in generative 3D Gaussians.
@@
Our method is also related to the Atlas representation initially proposed by AtlasNet [25] and its subsequent extensions [16, 55, 19]. AtlasNet models the shape as a union of independent MLPs, thus limiting the number of patches to a few dozen. In contrast, our proposed Atlas Gaussians model each patch using a patch center and patch features. This efficient encoding allows us to generate a significantly larger number of patches, providing stronger representation capabilities. Additionally, we use a transformer to learn this representation instead of an MLP, resulting in better scalability.
@@
Diffusion models. Diffusion models [90, 92, 91] have been dominant for diverse generation tasks, including image [33, 115, 75, 82], video [34, 31, 4], audio [37, 47, 54], text [51, 22, 53], and motion [116, 98, 111]. The success of these models typically follows the VAE + LDM paradigm. While pioneering efforts [112, 120, 63, 60, 118, 114, 13, 50, 43, 48] have attempted to apply this paradigm to 3D, the problem remains unsolved and has not achieved the same level of success. We argue that one of the main reasons is the need for an efficient VAE to represent high-quality 3D content, which is the key contribution of this paper.
@@
3D generation. 3D generation methods can be classified into two genres. The first is optimization-based methods [40, 93, 99, 52, 101, 9], which are time-consuming due to per-shape optimization. For example, DreamField [40] uses CLIP guidance. DreamFusion [76] and SJC [99] introduce 2D diffusion priors in different formats. Magic3D [52] employs a coarse-to-fine pipeline to improve convergence speed. Prolificdreamer [101] uses variational score distillation to enhance generation fidelity. Fantasia3D [9] disentangles geometry and texture to achieve higher quality generation.
@@
The second genre involves training a generalizable feed-forward network to output 3D content, which allows for fast 3D generation but often with less details. One direction is to apply existing generative modeling techniques (VAE [46], GAN [23], normalizing flow [81], autoregressive model [3, 24] and diffusion model) directly on various 3D representations. Another direction [106, 57, 56, 58, 84, 35, 97] uses 2D as intermediate representation, integrating 2D diffusion models [31, 86]. For example, Instant3D [49] trains a diffusion model to generate sparse multi-view images and uses a reconstruction model to derive the 3D shapes from the 2D images. VFusion3D [28] and V3D [12] use video diffusion models to improve the consistency between generated images. However, these methods rely on 2D representations and often suffer from multi-view consistency issues.
In this paper, we push the state-of-the-art in the first direction by developing a novel explicit representation Atlas Gaussians. Our representation is efficient and allows for learning 3D generative models from images.
@@ 3 Method
In this section, we first introduce our Atlas Gaussians representation (Sec. 3.1). Then we introduce how we learn a variational autoencoder to connect the 3D space with a latent space (Sec. 3.2). Finally, we introduce how we learn the generation model using latent diffusion in the learned latent space (Sec. 3.3).
@@ 3.1 Atlas Gaussians Representation
In 3D Gaussian Splatting [45], each 3D Gaussian $\boldsymbol{\mathrm{g}}$ can be parameterized with a center position $\boldsymbol{\mathrm{\mu}}\in\mathbb{R}^{3}$, scale $\boldsymbol{\mathrm{s}}\in\mathbb{R}^{3}$, rotation quaternion $\boldsymbol{\mathrm{r}}\in\mathbb{R}^{4}$, opacity ${\mathrm{o}}\in\mathbb{R}$ and color $\boldsymbol{\mathrm{c}}\in\mathbb{R}^{3}$. To achieve high-quality rendering results, typically a sufficiently large number of 3D Gaussians are required.
The key idea of Atlas Gaussians is to represent the shape as a union of $M$ local patches $\mathcal{A}=\{\boldsymbol{a}_{i}\}_{i=1}^{M}$, where each patch $\boldsymbol{a}_{i}$ can decode 3D Gaussians through UV-based sampling. As shown in Figure 2, we parameterize each local patch $\boldsymbol{a}_{i}:=(\boldsymbol{x}_{i},\boldsymbol{f}_{i},\boldsymbol{h}_{i})$ with patch center $\boldsymbol{x}_{i}\in\mathbb{R}^{3}$, the geometry features $\boldsymbol{f}_{i}\in\mathbb{R}^{4\times d}$ and the appearance features $\boldsymbol{h}_{i}\in\mathbb{R}^{4\times d}$. More specifically, we parameterize the geometry and appearance features as the features at the four corners of the local patch in the UV space. We denote $\boldsymbol{f}_{i}=(\boldsymbol{f}_{i1},\boldsymbol{f}_{i2},\boldsymbol{f}_{i3},\boldsymbol{f}_{i4})$ and $\boldsymbol{h}_{i}=(\boldsymbol{h}_{i1},\boldsymbol{h}_{i2},\boldsymbol{h}_{i3},\boldsymbol{h}_{i4})$. This geometry and appearance disentanglement can facilitate better learning [21, 121, 105]. We use the feature $\boldsymbol{f}_{i}$ to decode Gaussian positions while using $\boldsymbol{h}_{i}$ to decode the rest of Gaussian attributes. We also assign a 2D coordinate $\boldsymbol{u}_{ij}\in\mathbb{R}^{2}$ to each feature vector $\boldsymbol{f}_{ij}$ and $\boldsymbol{h}_{ij}$ as a positional embedding, where $\boldsymbol{u}_{i1}=(0,0)$, $\boldsymbol{u}_{i2}=(1,0)$, $\boldsymbol{u}_{i3}=(1,1)$, $\boldsymbol{u}_{i4}=(0,1)$, $\forall i$.
@@
For generation, we randomly sample query points in the UV space and each point will be decoded as a 3D Gaussian. Given a query point $\boldsymbol{q}_{ij}\in{[0,1]}^{2}$, we map the 2D coordinate $\boldsymbol{q}_{ij}$ to the center of 3D Gaussians as:
@@
where $\phi$ is the mapping function, which takes the query point location $\boldsymbol{q}_{ij}$, the pre-defined location of patch feature vectors $\boldsymbol{u}_{ij}$, and the geometry features $\boldsymbol{f}_{i}$ as inputs. We implement the mapping function using interpolation in the 2D space:
@@
One design choice for $w$ is the bilinear interpolation weight function, which has been widely used in feature decoding [70]. However, these linear weights purely based on coordinates have limited representation ability. Inspired by [114], we design a more powerful weight function defined on both coordinates and features:
@@
The benefits of Atlas Gaussians representation are three-fold. First, through UV-based sampling in a unit square, Atlas Gaussians enable easy generation of a sufficiently large number of 3D Gaussians. They also possess the potential to generate a variable and theoretically infinite number of 3D Gaussians. Second, Atlas Gaussians utilize a non-linear learnable weight function based on the MLP projection of the positional encoding, which has a stronger representation ability than the existing linear interpolation weight. Third, Atlas Gaussians are computation-efficient with low memory overhead. An important property of Atlas Gaussians is that they do not require extra network parameters when scaling up the number of generated 3D Gaussians. We also provide an ablation study in Section 4.4 to validate the nice properties of Atlas Gaussians.
@@
Figure 3: The proposed VAE architecture. CA denotes the cross-attention layer. The latent space defined by $\boldsymbol{z}_{0}$ is used for latent diffusion.

### v3

@@ Abstract
Using the latent diffusion model has proven effective in developing novel 3D generation techniques. To harness the latent diffusion model, a key challenge is designing a high-fidelity and efficient representation that links the latent space and the 3D space. In this paper, we introduce Atlas Gaussians, a novel representation for feed-forward native 3D generation. Atlas Gaussians represent a shape as the union of local patches, and each patch can decode 3D Gaussians. We parameterize a patch as a sequence of feature vectors and design a learnable function to decode 3D Gaussians from the feature vectors. In this process, we incorporate UV-based sampling, enabling the generation of a sufficiently large, and theoretically infinite, number of 3D Gaussian points. The large amount of 3D Gaussians enables the generation of high-quality details. Moreover, due to local awareness of the representation, the transformer-based decoding procedure operates on a patch level, ensuring efficiency. We train a variational autoencoder to learn the Atlas Gaussians representation, and then apply a latent diffusion model on its latent space for learning 3D Generation. Experiments show that our approach outperforms the prior arts of feed-forward native 3D generation. Project page: https://yanghtr.github.io/projects/atlas_gaussians.
footnotetext: * Equal Contribution
Figure 1: We propose the Atlas Gaussians representation for 3D generation. Our method supports both unconditional (left) and conditional (middle) generation with great diversity. With Atlas Gaussians, we can generate a sufficiently large, and theoretically infinite, number of 3D Gaussian points. To demonstrate this, 100K Gaussian points generated by our method are shown (right). Additionally, we iteratively sample 32 patches of the shape, displaying each set in one of four sub-figures (right).
@@ 1 Introduction
3D generation has become increasingly important in various domains, including virtual reality, gaming, and film production. Recent advances in diffusion models Ho et al. (2020); Song et al. (2020) have improved the quality of 3D generation, offering superior performance over previous methods, such as variational autoencoders (VAEs) Kingma & Welling (2013); Mo et al. (2019) and generative adversarial networks (GANs)Goodfellow et al. (2014); Gao et al. (2022).
@@
Despite progress, the effectiveness of 3D generation still falls short compared to 2D generation models. The robust performance in 2D generation is largely due to the effective integration of VAEs with latent diffusion models (LDMs) Rombach et al. (2022). A primary challenge hindering the complete success of this paradigm in 3D generation is the development of a high-fidelity 3D representation that can be efficiently embedded into a low-dimensional latent space. Pioneering efforts have applied diffusion models to various traditional 3D representations, such as point clouds Zeng et al. (2022); Zhou et al. (2021); Luo & Hu (2021), meshes Liu et al. (2023e), occupancy fields Zheng et al. (2023); Zhang et al. (2023a) and signed distance functions Cheng et al. (2023); Li et al. (2023b); Zhang et al. (2024). However, these approaches often focus solely on modeling geometry without considering the appearance attributes. More recently, a notable attempt Lan et al. (2024) has designed a VAE that uses volume rendering techniques Mildenhall et al. (2020) to incorporate appearance modeling. However, volume rendering presents inherent limitations, including slow rendering speeds and constrained rendering resolutions. To overcome these limitations, we have developed a VAE that leverages the latest 3D Gaussian representation Kerbl et al. (2023), which significantly improves both the quality and speed of rendering.
@@
Designing a VAE based on 3D Gaussians presents considerable challenges. The first challenge lies in creating an efficient decoder capable of mapping low-dimensional latents to 3D Gaussians. Existing methods for 3D Gaussian decoding focus predominantly on reconstruction tasks and typically lack an information bottleneck design Tang et al. (2024); Yinghao et al. (2024), thus failing to provide low-dimensional latent directly. Alternatively, they often require multiple complex and interdependent components Xu et al. (2024a); Zou et al. (2023). The second challenge involves generating a sufficiently large number of 3D Gaussians efficiently, since high-quality rendering necessitates an adequate quantity of these Gaussians. Some current methods Xu et al. (2024a); Zou et al. (2023) address this by employing additional complex point-upsampling networks to increase the number of 3D Gaussians, which inherently limits the number of Gaussians that can be generated. Other techniques Tang et al. (2024); Yinghao et al. (2024) utilize image representations to generate a large number of 3D Gaussians. However, for all these methods, more network parameters are usually required as the number of 3D Gaussians increases.
@@
To address the challenges of designing VAEs for 3D Gaussians, we propose Atlas Gaussians, a new representation for 3D generation. This representation is inspired by surface parameterization Floater & Hormann (2005), a foundation technique in many graphics applications where surface attributes are sampled and stored in a 2D texture map. Specifically, Atlas Gaussians model the shape as a union of local patches, with each patch decoding 3D Gaussians via UV-based sampling. By parameterizing 3D Gaussians in the UV space, we can easily generate a sufficiently large, and theoretically infinite, number of 3D Gaussians. Unlike traditional surface parameterization approaches, the UV mapping in Atlas Gaussians is learned end-to-end. A significant advantage of Atlas Gaussians is that the sampling process does not require additional network parameters as the number of 3D Gaussians increases.
@@
We design a transformer-based decoder to map low-dimensional latents to Atlas Gaussians. This decoder is specifically structured to disentangle geometry and appearance features, facilitating faster convergence and improved representation capabilities. Using the local awareness of Atlas Gaussians, we also reduce computational complexity by decomposing the self-attention layers. Finally, the latent space learned by our VAE can be applied to existing latent diffusion models efficiently.
@@
Note that in contrast to the main approach in 3D generation that uses the multi-view representation Wang & Shi (2023); Long et al. (2023); Liu et al. (2023d); Shi et al. (2023b), our approach is inherently 3D-based. Therefore, a key advantage is that we do not need to address the challenging multi-view consistency issue associated with the multi-view representation. Moreover, the rendering module of Atlas Gaussians allows representation learning from images.
@@
• We design a new transformer-based decoder to efficiently map low-dimensional latents to Atlas Gaussians, using separate branches to disentangle geometry and appearance features.
@@ 2 Related work
3D representation. 3D reconstruction and generation benefit from different 3D representations by leveraging their unique properties. These representations include explicit representations Wu et al. (2016); Mittal et al. (2022); Ren et al. (2024); Zeng et al. (2022); Zhou et al. (2021); Luo & Hu (2021); Sun et al. (2020a); Xie et al. (2021); Yang et al. (2019); Sun et al. (2020b); Achlioptas et al. (2018); Fan et al. (2017); Liu et al. (2023e); Nash et al. (2020); Siddiqui et al. (2023); Groueix et al. (2018); Chen et al. (2020); Kerbl et al. (2023); Tang et al. (2024); Yinghao et al. (2024); Zou et al. (2023); Szymanowicz et al. (2024) and implicit representations Park et al. (2019); Chen & Zhang (2019); Mescheder et al. (2019); Li et al. (2023b); Zheng et al. (2022); Shue et al. (2023); Jiang et al. (2022); Hui et al. (2022); Yan et al. (2022); Mildenhall et al. (2020); Zhang et al. (2022); Chan et al. (2022); Gu et al. (2023); Chen et al. (2023a); Cao et al. (2024); Müller et al. (2023); Watson et al. (2023). In this paper, we focus on 3D Gaussians Kerbl et al. (2023); Tang et al. (2024); Yinghao et al. (2024); Zou et al. (2023); Szymanowicz et al. (2024); He et al. (2024), which possess high-quality rendering procedures that allow learning from image supervisions. However, existing results mainly focus on reconstructing 3D Gaussians. Our goal is to push the state-of-the-art in generative 3D Gaussians.
@@
Our method is also related to the Atlas representation initially proposed by AtlasNet Groueix et al. (2018) and its subsequent extensions Deprelle et al. (2019); Liu et al. (2019); Feng et al. (2022). AtlasNet models the shape as a union of independent MLPs, thus limiting the number of patches to a few dozen. In contrast, our proposed Atlas Gaussians model each patch using a patch center and patch features. This efficient encoding allows us to generate a significantly larger number of patches, providing stronger representation capabilities. Additionally, we use a transformer to learn this representation instead of an MLP, resulting in better scalability.
@@
Diffusion models. Diffusion models Sohl-Dickstein et al. (2015); Song et al. (2021; 2020) have been dominant for diverse generation tasks, including image Ho & Salimans (2022); Zhang et al. (2023b); Podell et al. (2023); Rombach et al. (2022), video Ho et al. (2022b; a); Blattmann et al. (2023), audio Huang et al. (2023); Kong et al. (2020); Liu et al. (2023a) and text Li et al. (2022); Gong et al. (2022); Lin et al. (2023b). The success of these models typically follows the VAE + LDM paradigm. Although pioneering efforts Zeng et al. (2022); Zhou et al. (2021); Luo & Hu (2021); Liu et al. (2023e); Zheng et al. (2023); Zhang et al. (2023a); Cheng et al. (2023); Li et al. (2023b); Jun & Nichol (2023) have attempted to apply this paradigm to 3D, the problem remains unsolved and has not achieved the same level of success. We argue that one of the main reasons is the need for an efficient VAE to represent high-quality 3D content, which is the key contribution of this paper.
@@
3D generation. 3D generation methods can be classified into two genres. The first is optimization-based methods Jain et al. (2022); Sun et al. (2023); Wang et al. (2023); Lin et al. (2023a); Wang et al. (2024); Chen et al. (2023b), which are time-consuming due to per-shape optimization. For example, DreamField Jain et al. (2022) uses CLIP guidance. DreamFusion Poole et al. (2022) and SJC Wang et al. (2023) introduce 2D diffusion priors in different formats. Magic3D Lin et al. (2023a) employs a coarse-to-fine pipeline to improve convergence speed. Prolificdreamer Wang et al. (2024) uses variational score distillation to enhance generation fidelity. Fantasia3D Chen et al. (2023b) disentangles geometry and texture to achieve higher quality generation.
@@
The second genre involves training a generalizable feed-forward network to output 3D content, which allows for fast 3D generation but often with less details. One direction is to apply existing generative modeling techniques (VAE, GAN, normalizing flow Rezende & Mohamed (2015), autoregressive model Bengio et al. (2000); Graves (2013) and diffusion model) directly on various 3D representations. Another direction Xu et al. (2024b); Liu et al. (2024; 2023b; 2023c); Shi et al. (2023a); Hong et al. (2024); Tang et al. (2024) uses 2D as intermediate representation, integrating 2D diffusion models Ho et al. (2022a); Shi et al. (2023b). For example, Instant3D Li et al. (2023a) trains a diffusion model to generate sparse multi-view images and uses a reconstruction model to derive the 3D shapes from the 2D images. VFusion3D Han et al. (2024) and V3D Chen et al. (2024) use video diffusion models to improve consistency between generated images. However, these methods rely on 2D representations and often suffer from multi-view consistency issues. In this paper, we push the state-of-the-art in the first direction by developing a novel representation of Atlas Gaussians. Our representation is efficient and allows for learning 3D generative models from images.
@@ 3 Method
In this section, we first introduce our Atlas Gaussians representation (Sec. 3.1). Then we introduce how we learn a VAE to connect the 3D space with the latent space (Sec. 3.2). Finally, we introduce how we learn the generative model using latent diffusion in the learned latent space (Sec. 3.3).
@@ 3.1 Atlas Gaussians Representation
In 3D Gaussian Splatting Kerbl et al. (2023), each 3D Gaussian $\boldsymbol{\mathrm{g}}$ can be parameterized with a center $\boldsymbol{\mathrm{\mu}}\in\mathbb{R}^{3}$, scale $\boldsymbol{\mathrm{s}}\in\mathbb{R}^{3}$, rotation quaternion $\boldsymbol{\mathrm{r}}\in\mathbb{R}^{4}$, opacity ${\mathrm{o}}\in\mathbb{R}$ and color $\boldsymbol{\mathrm{c}}\in\mathbb{R}^{3}$. To achieve high-quality rendering results, typically a sufficiently large number of 3D Gaussians are required.
The key idea of Atlas Gaussians is to represent the shape as a union of $M$ local patches $\mathcal{A}=\{\boldsymbol{a}_{i}\}_{i=1}^{M}$, where each patch $\boldsymbol{a}_{i}$ can decode 3D Gaussians through UV-based sampling. As shown in Figure 2, we parameterize each local patch $\boldsymbol{a}_{i}:=(\boldsymbol{x}_{i},\boldsymbol{f}_{i},\boldsymbol{h}_{i})$ with patch center $\boldsymbol{x}_{i}\in\mathbb{R}^{3}$, the geometry features $\boldsymbol{f}_{i}\in\mathbb{R}^{4\times d}$ and the appearance features $\boldsymbol{h}_{i}\in\mathbb{R}^{4\times d}$. More specifically, we parameterize the geometry and appearance features as the features at the four corners of the local patch in the UV space. We denote $\boldsymbol{f}_{i}=(\boldsymbol{f}_{i1},\boldsymbol{f}_{i2},\boldsymbol{f}_{i3},\boldsymbol{f}_{i4})$ and $\boldsymbol{h}_{i}=(\boldsymbol{h}_{i1},\boldsymbol{h}_{i2},\boldsymbol{h}_{i3},\boldsymbol{h}_{i4})$. This type of feature disentanglement can facilitate more effective learning Gao et al. (2022); Zou et al. (2023); Xu et al. (2024a). We use the feature $\boldsymbol{f}_{i}$ to decode Gaussian positions while using $\boldsymbol{h}_{i}$ to decode the rest of Gaussian attributes. We also assign a 2D coordinate $\boldsymbol{u}_{ij}\in\mathbb{R}^{2}$ to each feature vector $\boldsymbol{f}_{ij}$ and $\boldsymbol{h}_{ij}$ as a positional embedding, where $\boldsymbol{u}_{i1}=(0,0)$, $\boldsymbol{u}_{i2}=(1,0)$, $\boldsymbol{u}_{i3}=(1,1)$, $\boldsymbol{u}_{i4}=(0,1)$, $\forall i$.
@@
For generation, we randomly sample query points in the predefined unit square UV space for each patch $\boldsymbol{a}_{i}$. Each point is then decoded into a 3D Gaussian. Given a query point $\boldsymbol{q}_{ij}\in{[0,1]}^{2}$, we map the 2D coordinate $\boldsymbol{q}_{ij}$ to the center of 3D Gaussians as:
@@
where $\phi$ is the mapping function, which takes the query point location $\boldsymbol{q}_{ij}$, the predefined location of patch feature vectors $\boldsymbol{u}_{i}$, and the geometry features $\boldsymbol{f}_{i}$ as inputs. We implement the mapping function using interpolation in the 2D space:
@@
One design choice for $w$ is the bilinear interpolation weight function, which has been widely used in feature decoding Müller et al. (2022). However, these linear weights purely based on coordinates have limited representation ability. Inspired by Zhang et al. (2023a), we design a more powerful weight function defined on both coordinates and features:
@@
The benefits of Atlas Gaussians representation are three-fold. First, through UV-based sampling in a unit square, Atlas Gaussians enable easy generation of a sufficiently large number of 3D Gaussians. They also possess the potential to generate a variable and theoretically infinite number of 3D Gaussians. Second, Atlas Gaussians utilize a non-linear learnable weight function based on the MLP projection of the positional encoding, which has a stronger representation ability than the existing linear interpolation weight. Third, Atlas Gaussians are computation efficient with low memory overhead. Another important property of Atlas Gaussians is that they do not require extra network parameters when scaling up the number of generated 3D Gaussians. We also provide an ablation study in Section 4.4 to validate the nice properties of Atlas Gaussians.
@@
Figure 3: The proposed VAE architecture. CA denotes the cross-attention layer. For simplicity, the variational component of the VAE is omitted. The latent $\boldsymbol{z}_{0}$ is used for latent diffusion.